# From Cash to Cashless: UPI’s Impact on Spending Behavior among Indian Users

### v2

@@ Abstract.
The emergence of digital payment systems has transformed how individuals conduct financial transactions, offering convenience, security, and efficiency. One groundbreaking innovation making waves in the Indian financial landscape is the Unified Payments Interface (UPI). Existing work has explored how digital payments benefit a country’s economy and GDP. However, our study explores how the introduction of UPI has influenced spending behavior among Indian users on an ”individual” level. We gathered 235 valid survey responses encompassing diverse demographics and interviewed 20 survey respondents. Approximately 75% of the survey respondents reported increased spending due to UPI, with only 7% indicating reduced spending. Significantly, 91.5% of the respondents reported satisfaction with their UPI usage. Also, 95.2% of the survey respondents found making payments via UPI convenient. Our research also provides suggestions for UPI applications and various stakeholders to enhance digital payment systems, enabling users to make informed decisions and fostering responsible financial management. *Equal Contribution
Digital Payments, Cashless Transactions, Unified Payments Interface (UPI), User Study
journalyear: 2024copyright: rightsretainedconference: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems (CHI EA ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613905.3651050isbn: 979-8-4007-0331-7/24/05ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
An E-payment is generally understood as an economic situation where the buying and selling of goods are handled without physical cash, often facilitated through E-payment or checks (Pau, 2012). Digital payments have become integral to modern financial landscapes, offering swift, secure, and convenient electronic transactions that transcend traditional methods. This global shift towards digital transactions has been accelerated by the widespread adoption of internet and mobile technologies, reshaping financial practices across the world (Alkhowaiter, 2020). In 2022 alone, the volume of real-time digital transactions reached an unprecedented 195 million (89., 2023).
@@
Despite UPI’s monumental success, a critical research gap exists regarding its influence on users’ spending habits. Our study addresses this crucial gap in existing research on UPI, specifically focusing on the impact of UPI on users’ spending habits. While previous studies and works (Sreenu, 2020) acknowledge UPI’s success in reshaping digital payments, they fall short of providing detailed insights into how this technology influences ”individual” spending behaviors. The existing literature often emphasizes macro-level implications and adoption rates. Multiple questions about changes in expenditure preferences, budgeting strategies, investment choices, long-term financial goals and overall financial behaviors remain unexplored, hindering a holistic understanding of UPI adoption’s intended and unintended consequences on users’ daily financial practices. This study seeks to fill these critical gaps by providing in-depth insights into the micro-level effects of UPI on spending behaviors. Moreover, there’s a notable absence of research delving into potential variations in UPI’s impact across different demographic groups like age, profession, and socio-economic status, etc.
@@
We explore users’ perceptions of security and how trust in UPI influences expenditure patterns, unraveling the intricate dynamics between trust, technology adoption, and financial decision-making. Additionally, we investigate the relationship between users’ financial literacy levels and UPI utilization, shedding light on how understanding financial principles shapes spending choices. Our research also analyzes the existing budgeting applications, employing a competitive lens to understand its impact on users’ awareness of spending patterns and financial discipline.
The significance of this study extends beyond the realm of finance. In Human-Computer Interaction (HCI), understanding how users interact with financial technologies is essential for designing user-friendly and effective interfaces. This study contributes valuable insights to HCI researchers, financial institutions, and policymakers, guiding the development of digital payment systems that align with users’ financial needs and behaviors.
In addressing the above-mentioned gaps, this research employs a mixed-methods approach, encompassing qualitative and quantitative assessments. We gathered 276 responses covering diverse demographics and backgrounds using a survey form, out of which 235 were valid responses. Additionally, we also conducted semi-structured interviews with 20 survey respondents. This study aims to address the following research questions (RQs):
@@
    RQ3: To what extent does UPI adoption vary across demographic groups, such as age and profession, and how do these variations influence users’ financial practices, including spending habits and long term financial goals?
Among the participants we surveyed, 74.2% of the survey participants confirmed increased spending due to UPI, with only 7% confirming reduced spending. Significantly, 91.5% of the survey respondents reported satisfaction with their UPI usage and 95.2% of the survey respondents found making payments via UPI convenient. Our research also provides suggestions for UPI applications and various stakeholders to enhance digital payment systems, enabling users to make informed decisions and fostering responsible financial management. To the best of our knowledge, this is the first user study of UPI’s impact on individual spending behaviour in India.
@@ 2. Related Work
3. Methodology
3.1. RESEARCH DESIGN
@@
We employed a mixed methods (mul, [n. d.]) approach to investigate the impact of UPI on individual spending behaviors and financial practices. We utilized two primary data collection techniques: surveys and interviews. We used a random sampling method for the survey component, targeting individuals aged 18 and above. Our survey was conducted through Google Forms to ensure easy accessibility and respondent anonymity (by not asking for their personal information like name, phone number, etc). It garnered 276 responses from a diverse participant pool spanning various age groups and occupations. To ensure the quality, consistency and relevance of the responses, we implemented stringent criteria for validity. Frivolous responses containing irrelevant information were excluded, for instance, in the question regarding the initiation of UPI usage, any response indicating usage before 2016 was deemed invalid, considering that UPI was publicly launched that year. This validation process resulted in 235 valid responses. We conducted in-depth semi-structured interviews by randomly selecting interviewees from our survey respondents who expressed an interest in participating. We conducted 20 (13 male & 7 female) interviews and they provided verbal consent for the same. Survey and interview questions are presented in Appendix A and Appendix B respectively.
@@ 3.2. ETHICAL CONSIDERATIONS
In conducting this research, we implemented various ethical considerations to uphold transparency and safeguard the privacy and well-being of participants. Our study’s research materials and protocols underwent thorough review and approval by our university’s Institutional Review Board (IRB). Before participating in interviews, all candidates received information about the study’s purpose, the voluntary nature of their involvement, and the assurance of anonymity and confidentiality. We also acknowledge the significance of capturing diverse perspectives and experiences of individuals using UPI and its influence on their spending behavior. To ensure inclusivity, we actively sought participants from various backgrounds and genders. Verbal consent was obtained from participants for recording the interviews.
@@ 3.3. LIMITATIONS
One notable constraint of this study is the geographic scope of our data collection efforts. The research exclusively focused on respondents from urban areas in India, where technology adoption and usage, including electronic payment methods like UPI, tend to be higher and well integrated into daily life. As a result, the findings may not encompass the diverse spending habits. They may not accurately represent the impact of UPI on the spending habits of those living in rural areas in India.
4. Evaluations
4.1. QUANTITATIVE ANALYSIS OF SURVEY RESPONSES
@@
Our survey received 235 valid responses from a diverse range of respondents from various demographics and backgrounds. Among the respondents, 51.6% identified as students, 42% as working professionals, and 4.7% as business owners. Regarding gender distribution, the data shows that 77% of respondents were male, 19.7% were female, and 3.3% preferred not to disclose their gender. Regarding the adoption timeline of UPI (figure 1), we can conclude that the diverse range of adoption dates reflects users’ varying entry points into the UPI ecosystem, often aligning with significant events like demonetization and pandemic-related lockdowns.
@@
To understand the impact of UPI adoption, 74.2% of the survey respondents admitted that UPI had led to increased spending, while only 7% said that it had decreased spending. Additionally, 18.8% of the survey respondents felt no changes in their spending habits after adopting UPI. Regarding satisfaction levels (figure 3), 55% of the survey respondents indicated they were highly satisfied with their overall UPI experience. Additionally, 36.5% reported satisfaction, and 7.7% remained neutral in their assessment. Therefore we can conclude that 91.5% (55% + 36.5%) of the respondents are satisfied with their overall experience using UPI for payments. When asked to rate their convenience with UPI for making payments (figure 7), the survey results indicate that 69.50% of respondents rated ”Highly convenient”, 25.7% rated ”Convenient, and 3.7% were ”Neutral” on the same. Therefore we can conclude that 95.2% (69.5% + 25.7%) of the respondents find using UPI for payments convenient. Notably, only 2-3 respondents reported dissatisfaction with UPI as a payment method, indicating that most were satisfied with their experience.
@@
Respondents reported a varied daily spending range using UPI, from Rs 50 (0.60 USD) to 1 lakh (1,202.66 USD). On average, 15.7% mentioned a daily spending of around Rs 200 (2.41 USD). Regarding the frequency of UPI usage for payments (figure 2), a significant 81.1% use it daily, 15.6% use it weekly, 1.1% use it monthly, 1.90% use it occasionally, and 0.30% use it rarely. This data highlights the widespread and frequent adoption of UPI among surveyed individuals.
@@
When asked about overspending due to UPI usage (figure 5), 59.8% of respondents acknowledged experiencing such situations, while 39.8% believed they had not. Additionally, respondents had diverse opinions when asked about the impact of using UPI on their budget adherence (figure 4); approximately 23% strongly agreed, 23.3% agreed, 26.7% remained neutral, 15.6% disagreed, and 11.5% reported strongly disagree, indicating that UPI has had an impact on their ability to stick to their budget and control overspending.
@@
Respondents employed various methods to manage their expenses, like checking their transaction history, Splitwise (spl, 2011) SBI YONO (yon, 2017), Samsung Notes (sno, 2016), using excel sheets and dedicated finance apps like MyMoneyy (mym, 2022), Honeybook (hon, 2013), Monefy (my, 2016), Payhawk (pay, 2018) and Money Manager (mm, 2020). This observation highlights diverse financial management approaches among UPI users, showcasing varied responses and opinions. A significant proportion acknowledges increased spending along with high overall satisfaction with UPI and its convenience. Our comprehensive analysis of various financial tracking and budget management apps in the Indian market is summarized in Table 1 in Section 5.
Figure 1. When did users start using UPI?
Figure 2. How often do users use UPI?
Figure 3. Satisfaction Level with UPI
Figure 4. UPI Impact on Budget
Figure 5. Budget exceedance due to UPI
@@
Figure 6. How often do people keep track of transaction history?
@@
4.2. QUALITATIVE ANALYSIS OF INTERVIEWS

### v3

@@ Abstract.
Unified Payments Interface (UPI) is a groundbreaking innovation making waves in digital payment systems in India. It has revolutionised financial transactions by offering enhanced convenience and security. While previous research has primarily focused on the macroeconomic effects of digital payments, our study examines UPI’s impact on individual spending behavior. Through a survey of 276 respondents and 20 follow-up interviews, we found that approximately 75% of participants reported increased spending due to UPI. Many attributed this to UPI’s intangible nature, which reduced feelings of guilt typically associated with spending. Additionally, participants provided suggestions to improve the user experience of existing UPI applications. Utilizing this feedback, we developed a high-fidelity prototype based on a popular UPI app in India and conducted usability testing with 34 participants. The insights gathered from this testing shaped the final prototype and its features. This study offers valuable design recommendations for UPI app developers and other stakeholders.
Digital Payments, Cashless Transactions, Unified Payments Interface (UPI), User Study, Usability Testing, Technology Acceptance Model, Perceived Usefulness, Perceived Ease of Use
$c$s: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Digital payments have become integral to modern financial landscapes, offering swift, secure, and convenient electronic transactions that transcend traditional methods. This global shift towards digital transactions has been accelerated by the widespread adoption of internet and mobile technologies, reshaping financial practices across the world (Alkhowaiter, 2020). In 2022 alone, the volume of real-time digital transactions reached an unprecedented 195 million (89., 2023).
@@
Despite UPI’s monumental success, a critical research gap exists regarding its influence on users’ spending habits. Our study addresses this crucial gap in existing research on UPI, specifically focusing on analyzing the impact of UPI on users’ spending habits. While previous studies (Sreenu, 2020) acknowledge UPI’s success in reshaping digital payments, they fall short of providing detailed insights into how this technology influences ”individual” spending behaviors. The existing literature often emphasizes macro-level implications and adoption rates. Multiple questions about changes in expenditure preferences, budgeting strategies, and overall financial behaviors remain unexplored, hindering a holistic understanding of UPI adoption’s intended and unintended consequences on users’ daily financial practices. This study seeks to fill these critical gaps by providing in-depth insights into the micro-level effects of UPI on spending behaviors.
@@
Moreover, there’s a notable absence of research delving into potential variations in UPI’s impact on spending behaviour across different demographic groups like age, profession, socio-economic status, etc. Our study aims to bridge this gap as well. We also explore users’ perceptions of security and how trust in UPI influences expenditure patterns, unraveling the intricate dynamics between trust, technology adoption, and financial decision-making. Additionally, we investigate the relationship between users’ financial management levels and UPI utilization, shedding light on how understanding financial principles shapes spending choices.
Phase 1: The above mentioned portion of the study constitutes Phase 1. In this phase, we employed a comprehensive mixed-methods approach to address the identified gaps, integrating qualitative and quantitative analyses. We began by distributing an initial survey to a diverse demographic to assess UPI usage, which resulted in 276 total responses, with 235 being valid. To gain deeper insights into user behavior, pain points, and preferences, we conducted semi-structured interviews with 20 respondents from the survey.
Phase 2: Based on the feedback and recommendations gathered in Phase 1, we developed an initial prototype based on a popular UPI app in India (PhonePe (pho, 2015, 2023)) in Phase 2. We selected PhonePe as the base UPI app due to its significant 48.9% market share, making it the largest UPI platform used in India (See Figure 1). To finalize the requirements for the initial prototype, we also analyzed the existing budgeting applications, employing a competitive lens to understand their impact on users’ awareness of spending patterns and financial discipline. The developed prototype incorporated several new features, including an expense tracker, balance display, and budgeting limits, aimed at improving financial awareness and promoting responsible spending as detailed in §7.1.
We then conducted usability testing on the prototype, involving 1-on-1, multi-stage personalized sessions with 34 participants. These sessions aimed to validate the prototype and gather detailed user feedback on new features and overall user experience. The usability testing provided valuable insights into users’ preferences, areas of challenge, and further improvements. This iterative process led to the development of a refined prototype that not only met users’ needs but also incorporated the most-requested features, with a focus on promoting responsible financial behavior and enhancing overall usability.
Figure 1. MARKET SHARE OF UPI PAYMENT APPS IN INDIA
This study aims to address the following research questions (RQs):
@@
    RQ3 : To what extent does UPI adoption vary across demographic groups, such as age and profession, and how do these variations influence users’ financial practices, including spending habits and long-term financial goals?
  4. (4)
    RQ4 : What are the strengths and limitations of current UPI apps from a user perspective, and how can their design be refined to improve overall user satisfaction?
Among the participants in Phase 1, 74.2% reported increased spending due to UPI, while only 7% noted a decrease in spending. Notably, 91.5% of respondents expressed satisfaction with their UPI usage, and 95.2% found making payments via UPI to be convenient.
In Phase 2, based on the usability testing of the initial prototype, 87.9% of users indicated that adding an expense tracker to financial apps like PhonePe (pho, 2024) and Paytm (pay, 2010) would be beneficial. This would provide them the ability to view expenses by month and category (e.g., food, travel), with a day-wise breakdown and budget allocation for each component. Additionally, 78.7% of users agreed that seeing their balance before and after making a payment would be advantageous.
This study offers valuable insights for HCI researchers, financial institutions, and policymakers, providing guidance for the development of digital payment systems that align with users’ financial needs, behaviors, and promote responsible financial management. To the best of our knowledge, this is the first user study in India that explores UPI’s impact on individual spending behavior and prototypes UPI interfaces designed for financial responsibility.
@@ 2. Related Work
The use of UPI in India has been on the rise, but at the same time the number of UPI related scams have also been increasing. The massive growth of UPI, however, has has attracted the attention of fraudsters, who are increasingly targeting vulnerable users. According to Indian government figures, there were more than 95,000 cases of fraud involving UPI in the financial year ending April 2023, up from 77,000 in the previous year (sca, 2024). According to Dixit et al. (bus, 2024) , the reason why there is a surge in number of UPI scams is not just the growing popularity of digital payments, but also a lack of financial literacy and imprudent use of technology, rendering a vast population vulnerable to such attacks. For instance, Allums et al. (exp, 2023) underscores the importance of creating frictionless interactions and ensuring robust data protection to meet user expectations, particularly in payment scenarios . Surendran et al. (Surendran et al., 2018) focuses on the need for simplifying e-payment apps to improve adoption among users with limited technology access and low literacy, using apps like WhatsApp as examples of effective simplification . Gupta et al. (eld, 2024) emphasizes the development of secure e-wallet authentication designs, incorporating device-specific identity features and a two-phase authentication process to enhance security and reliability . Additionally, Hassan et al. (Hassan and Shukur, 2022)on mobile expense tracking applications aims to create user-friendly solutions for young adults, helping them manage and visualize their daily transactions and budgeting . Finally, Kiat et al. (cas, 2023) on prototype design and recommendations for e-wallet apps discusses enhancing inclusivity but does not delve deeply into the iterative design process or how user feedback integrates into refining prototypes.
Our research fills this gap enhances app functionality by grounding the prototype development in iterative testing and real-world feedback. It significantly boosts user satisfaction, ensuring the final product is practical and user-centric.
3. Overview of the Study
This section provides an overview of our study on the impact of UPI on spending behavior and the development of financially responsible UPI interfaces (see Figure 2).
Phase 1: In §4, we outline the methodology used to collect data on UPI’s influence on users’ spending patterns. §5 presents the results of both quantitative and qualitative analyses of the collected data. Finally, §6 summarizes the key insights from our data analysis and provides a competitive analysis of existing personal finance tracking apps.
Phase 2: Building on the insights from Phase 1 regarding UPI’s impact on spending behavior and identifying key gaps, Phase 2 focuses on prototype development and its usability testing. First, we developed the initial prototype (version 1) whose details are presented in §7. §8 presents the details of the usability testing of the developed prototype. The results of this usability study are presented in §9. In §10, we address the user feedback and recommendations gathered during usability testing, which informed the development of the refined final prototype (version 2). Based on these findings, §11 provides recommendations for UPI app developers and stakeholders, focusing on improving user engagement and enhancing financial awareness. Lastly, §12 revisits the research questions introduced in the beginning and offers answers based on the study’s findings. We conclude in §13.
Figure 2. Overview of the entire research study
4. Phase 1 : Methodology
4.1. Data Collection
@@
We employed a mixed-methods approach to investigate the impact of UPI on individual spending behaviors and financial practices. We utilized two primary data collection techniques: surveys and interviews. We used a random sampling method for the survey component, targeting individuals aged 18 and above. Our survey was conducted through Google Forms to ensure easy accessibility and respondent anonymity (by not asking for their personal information like name, phone number, etc). It garnered 276 responses from a diverse participant pool spanning various age groups and occupations. To ensure the quality, consistency and relevance of the responses, we implemented stringent criteria for validity. Frivolous responses containing irrelevant information were excluded, for instance, in the question regarding the initiation of UPI usage, any response indicating usage before 2016 was deemed invalid, considering that UPI was publicly launched that year. This validation process resulted in 235 valid responses. We conducted in-depth semi-structured interviews by randomly selecting interviewees from our survey respondents who expressed an interest in participating. We conducted 20 (13 male & 7 female) interviews after receiving explicit consent from each of the participants. Survey and interview questions are presented in Appendix A and Appendix B, respectively.
4.2. Data Analysis
The data collected in Phase 1 was analyzed using both quantitative and qualitative methods, reflecting the mixed-methods approach employed in this phase. For the quantitative analysis, we aggregated and analyzed the survey data to identify overall trends and patterns in UPI usage and its impact on spending behavior. Key metrics such as the percentage of users reporting increased or decreased spending, satisfaction levels with UPI, and perceptions of convenience were derived from the survey responses. These aggregate numbers highlight the broader behavioral trends among UPI users across different demographics.
For the qualitative analysis, we conducted a thematic analysis of the interview transcripts. Using an inductive approach, we identified recurring themes and patterns related to user experiences, perceptions, and financial practices. This method allowed us to extract meaningful insights directly from the data without imposing pre-determined categories. Themes such as spending habits, security concerns, budgeting strategies, and recommendations for improving UPI applications were coded and categorized. The detailed thematic analysis provided a deeper understanding of individual perspectives and behaviors, complementing the broader trends observed in the survey data.
@@ 4.3. Ethical Considerations
In conducting this research, we implemented various ethical considerations to uphold transparency and safeguard the privacy and well-being of participants. Our study’s research materials and protocols underwent thorough review and approval by our university’s Institutional Review Board (IRB). Before participating in this study, all candidates received information about the study’s purpose, the voluntary nature of their involvement, and the assurance of anonymity and confidentiality. Explicit consent was obtained from participants to record the interviews.
@@ 4.4. Limitations
The research exclusively focused on respondents from urban areas in India, where technology adoption and usage, including electronic payment methods like UPI, tend to be higher and well integrated into daily life. As a result, the findings may not encompass the diverse spending habits. They may not accurately represent the impact of UPI on the spending habits of those living in rural areas in India.
5. Phase 1: Evaluating the Impact of UPI on Spending Behaviour
5.1. Quantitative Analysis (Analysis of Survey Results)
@@
Our survey received 235 valid responses from a diverse range of respondents from various demographics and backgrounds. Among the respondents, 51.6% identified as students, 42% as working professionals, and 4.7% as business owners. Regarding gender distribution, 77% of respondents were male, 19.7% were female, and 3.3% preferred not to disclose their gender. Regarding the adoption timeline of UPI (figure 3(a)), we can conclude that the diverse range of adoption dates reflects users’ varying entry points into the UPI ecosystem, often aligning with significant events like demonetization and pandemic-related lockdowns.
@@
To understand the impact of UPI adoption, 74.2% of the survey respondents admitted that UPI had led to increased spending, while only 7% said that it had decreased spending. Additionally, 18.8% of the survey respondents felt no changes in their spending habits after adopting UPI. Regarding satisfaction levels (figure 4(a)), 55% of the survey respondents indicated they were highly satisfied with their overall UPI experience. Additionally, 36.5% reported satisfaction, and 7.7% remained neutral in their assessment. Therefore we can conclude that 91.5% (55% + 36.5%) of the respondents are satisfied with their overall experience using UPI for payments. When asked to rate their convenience with UPI for making payments (figure 6), the survey results indicate that 69.50% of respondents rated ”Highly convenient”, 25.7% rated ”Convenient, and 3.7% were ”Neutral” on the same. Therefore we can conclude that 95.2% (69.5% + 25.7%) of the respondents find using UPI for payments convenient. Notably, only 2-3 respondents reported dissatisfaction with UPI as a payment method, indicating that most were satisfied with their experience.
@@
Respondents reported a varied daily spending range using UPI, from Rs 50 (0.60 USD) to 1 lakh (1,202.66 USD). On average, 15.7% mentioned a daily spending of around Rs 200 (2.41 USD). Regarding the frequency of UPI usage for payments (figure 3(b)), a significant 81.1% use it daily, 15.6% use it weekly, 1.1% use it monthly, 1.90% use it occasionally, and 0.30% use it rarely. This data highlights the widespread and frequent adoption of UPI among surveyed individuals.
@@
When asked about overspending due to UPI usage (figure 5(a)), 59.8% of respondents acknowledged experiencing such situations, while 39.8% believed they had not. Additionally, respondents had diverse opinions when asked about the impact of using UPI on their budget adherence (figure 4(b)); approximately 23% strongly agreed, 23.3% agreed, 26.7% remained neutral, 15.6% disagreed, and 11.5% reported strongly disagree, indicating that UPI has had an impact on their ability to stick to their budget and control overspending.
@@
Respondents employed various methods to manage their expenses, like checking their transaction history, Splitwise (spl, 2011) SBI YONO (yon, 2017), Samsung Notes (sno, 2016), using excel sheets and dedicated finance apps like MyMoneyy (mym, 2022), Honeybook (hon, 2013), Monefy (my, 2016), Payhawk (pay, 2018) and Money Manager (mm, 2020). This observation highlights diverse financial management approaches among UPI users, showcasing varied responses and opinions. A significant proportion acknowledges increased spending along with high overall satisfaction with UPI and its convenience. Our comprehensive analysis of various financial tracking and budget management apps in the Indian market is summarized in Table 1 .
5.2. Graphs for Quantitative Results
Figure 3. (a) When did users start using UPI? (b) How often do users use UPI?
Figure 4. (a) Satisfaction Level with UPI (b) UPI Impact on Budget
@@ 5.2. Graphs for Quantitative Results
Figure 5. (a) Budget exceedance due to UPI (b) How often do people keep track of transaction history?
@@
5.3. Qualitative Analysis (Analysis of Interview Results)